Determine the upper-tail critical value ta/2 in each of the following circumstances. a. 1-a = 0.90, n = 65 b. 1 - a=0.95, n =ROUND ANSWER TO FOUR DECIMAL PLACES
Answer #1

a. 1-\small \alpha=0.90, n= 65 ;

\small \alpha=1-0.90 =0.10 ; \small \alpha /2 =0.10/2 =0.05 ; Degrees of freedom = n-1 = 65-1=64; t\small \alpha/2 = t0.05

For 64 degrees of freedom , t0.05 = 1.6690

Answer : 1.6690

b. 1-\small \alpha=0.95, n= 65 ;

\small \alpha=1-0.95 =0.05 ; \small \alpha /2 =0.05/2 =0.025 ; Degrees of freedom = n-1 = 65-1=64; t\small \alpha/2 = t0.025

For 64 degrees of freedom , t0.025 = 1.9977

Answer : 1.9977

c. 1-\small \alpha=0.90, n= 54 ;

\small \alpha=1-0.90 =0.10 ; \small \alpha /2 =0.10/2 =0.05 ; Degrees of freedom = n-1 = 54-1=53 ; t\small \alpha/2 = t0.05

For 53 degrees of freedom , t0.05 =

Answer : 1.6741

d. 1-\small \alpha=0.90, n= 26 ;

\small \alpha=1-0.90 =0.10 ; \small \alpha /2 =0.10/2 =0.05 ; Degrees of freedom = n-1 = 26-1=25 ; t\small \alpha/2 = t0.05

For 25 degrees of freedom , t0.05 = 1.7081

Answer : 1.7081

e. 1-\small \alpha=0.99, n= 36 ;

\small \alpha=1-0.99 =0.01 ; \small \alpha /2 =0.01/2 =0.005 ; Degrees of freedom = n-1 = 36-1=35 ; t\small \alpha/2 = t0.005

For 35 degrees of freedom , t0.005 = 2.7238

Answer : 2.7238